A large tank contains 100 gallons of pure water. How many gallons of a saline solution containing 25% salt must be added to obtain a solution that has 10% salt ?
Answer by mananth(16946)

water 100 gallons 0%
saline 25% x gallons
total 10% x+100
..
0+ 25x=10(x+100)
25x=10x+1000
25x-10x=1000
15x= 1000
x= 1000/15
x= 66.66 gallons of water